OpenStudy (anonymous):

Can someone tell me how to simplify radicals? its on my final today and my teacher didnt even teach it

OpenStudy (anonymous):

what kind of radicals

OpenStudy (anonymous):

like radical 60x4y2

OpenStudy (anonymous):

thats x to the 4th power

OpenStudy (anonymous):

\(\sqrt{60x^4y^2}\) note that 60 is \(2^2 * 3 * 5\)\[=\sqrt{2^2 * 3 * 5 * x^4 * y^2}\]\[=2x^2|y|\sqrt{15}\]

OpenStudy (anonymous):

wait why is it 2 to the 2nd power?

OpenStudy (anonymous):

the general rule is to look for parts of the expression that you can easily take OUT of the square root sign. for example, \[\sqrt{60}\] = \[\sqrt{4*15}\] =\[2\sqrt{15}\]

OpenStudy (anonymous):

2 to the 2nd power = 4 4*3*5 = 60

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

you have to break it down until you can take things out of the square root sign, like an X^2 can be taken out of the square root sign and turned into an X.

OpenStudy (anonymous):

does that make sense aavarino?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you even do the square root of the exponents

OpenStudy (anonymous):

im learning this on my own my teacher didnt even show us how to do these lol

OpenStudy (anonymous):

and to simplify should all the variables be outside of the radical?

OpenStudy (anonymous):

yup! exactly. \[\sqrt{x ^{4}}=x ^{2}\]

OpenStudy (anonymous):

the goal of simplifying is to get AS MUCH AS YOU CAN out of the radical sign.

OpenStudy (anonymous):

\[\sqrt{y ^{2}}=y\]

OpenStudy (anonymous):

it's possible that a variable (x or y) may be left inside the radical sign, and that's okay, as long as it's as simple as you can make it.

OpenStudy (anonymous):

\[\sqrt{27x ^{4}y}\]

OpenStudy (anonymous):

for that one is it \[3x ^{2}y \sqrt{3} ? \]

OpenStudy (anonymous):

or did i do it wrong lol

OpenStudy (anonymous):

ALMOST, you're very close. the y that is under the radical - you can't reduce that any further. so it stays under the radical along with the 3 you already have under there.

OpenStudy (anonymous):

oh!! i gotcha so whatever is already simplified leave in the radical

OpenStudy (anonymous):

IF IF IF it was a y^2 under the radical, then you can reduce that to just y, but it isn't in this probelm.

OpenStudy (anonymous):

yes, if it's in there to begin with, it must stay in there. it's like a box. you want to take out the stuff you can, and what's left inside is already as simple as it can get. like a 3, or a y. if you have a 4 left in there though ,then you have to take it out (square root of 4 = 2).

OpenStudy (anonymous):

can i give you another one to see if you can do it now?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

this one is hard, but i think you can do it: \[\sqrt{40x ^{3}y ^{4}}\]

OpenStudy (anonymous):

wait quick question before i do it.. since 40 has a lot of factors can i choose any? would i get the same answer if i chose 10 and 4 or 8 and 5?

OpenStudy (anonymous):

great question! sure, just remember that what's left in the radical has to be as simple as possible. so let's say you break 40 down into: 8 and 5. MAKE SURE YOU NOTICE that 8 can then be further broken down into 4 and 2. so really, 40 can be broken down into 4 and 2 and 5. OR 40 can be broken down into 4 and 10. try it both ways, and see if you get the same answer. :-)

OpenStudy (anonymous):

\[2y ^{2} \sqrt{4x ^{3}} ?\]

OpenStudy (anonymous):

oops i lied

OpenStudy (anonymous):

i meant 2 in the radical lol

OpenStudy (anonymous):

let's look at just the X^3 under the radical for a second: \[\sqrt{x ^{3}} = \sqrt{x ^{2}x}\] how would you reduce that part?

OpenStudy (anonymous):

i thought 3 didnt reduce? but the x^2 reduces to x

OpenStudy (anonymous):

the second part is right, but X^3 is the same as X*X*X = X^2*X all those expressions are the same.

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so how would the answer look from the question before?

OpenStudy (anonymous):

you had some part of it right: \[2xy ^{2}\sqrt{10x}\] make sure you can get to that, and if not, ask me for more help please :-)

OpenStudy (anonymous):

question.. but cant 10 be simplified?

OpenStudy (anonymous):

it can be broken down into 5 and 2, however neither of those is a clean square root so it doesn't make it any simpler. (good question!)

OpenStudy (anonymous):

oh ok! for not learning this in class i understand now

OpenStudy (anonymous):

well getting the hang of it

OpenStudy (anonymous):

great! should we try another one?

OpenStudy (anonymous):

i have 2 more on my review sheet. \[\sqrt{100x}/121 \] ( not sure how to do a fraction on here )

OpenStudy (anonymous):

\[\sqrt{100x}/121\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but its all in the radical

OpenStudy (anonymous):

oops, that's what i was going for :-)

OpenStudy (anonymous):

okay, so can either 100, x, or 121 be reduced?

OpenStudy (anonymous):

yes 10, and 11

OpenStudy (anonymous):

awesome. but as you noticed the X has to stay inside the radical.

OpenStudy (anonymous):

so what would you get?

OpenStudy (anonymous):

radical 10x/11 ? does the radical go over the whole fraction?

OpenStudy (anonymous):

note that \[\sqrt{100}= 10 the 10 is OUTSIDE the radical. same for the 11.

OpenStudy (anonymous):

so the x is the only thing that stays in the radical?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

so you should have